- 15 Hz ops. TLG changes to do 5 hours continuous at 15 Hz. Demonstrated for future.
- recycler kickers have long rise time, need 3.
- our trigger is set from first bunch filled, so not an issue. We would see one less bunch per spill if we were sensitive.
- April 6, measurements on cable lengths.
- 3 bunch notch from a 2 bunch notch
- fast trim magnent is identified and tested
- hardening on seal for horn 4 being tested if it’s just seal, or throughout bulk
- Air fan replaced in target hall. (Backfish)
- Dedicated collimator studies next week.
- Last week commissioning HV system.
- Fill tank w/ water this week Wednesday, ~8 hours.
- Downtime last week: test noise filter on service board Wed-Thur.
- Factor x2 reduction in noise.
- Study compression ratios of data & implications for physics.
- Issues w/ Sunday night SlowMon computer: RAID array firmware is updated. Issue went away last week.
- Full filter used instead of reduced one because reduced one was not filtering enough oxygen from the top-off argon.
- Spare pump put into use, old pump was used for ~20% longer w/out maintenance.
- Have weird # of triggers per spill